OpenStudy (anonymous):

plz help The radius of a cylinder is 3.5 ft. The height is 14 ft. Find the surface area and volume of the cylinder to the nearest tenth of a foot. Show your work.

OpenStudy (anonymous):

the main thing to know is that all a cylindar is is two circles and a rectangle bent into a circle. you know the formula for a circle is A=pi*r^2 the only other part is the rectangle. You know the height is 14 so what's the width

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

the width is the circumference of the circle, C = 2*pi*r

OpenStudy (anonymous):

so what you have is 2 circles and a rectangle 2*pi*r^2 + h*(2*pi*r) = surface area

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

hopw that makes some sense

OpenStudy (anonymous):

hope*

OpenStudy (anonymous):

thank

OpenStudy (anonymous):

you